Solve the differential equation:
dy/dx = e^(xy)

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear DEEPAK

dy/dx = e^(xy)

let xy=t

 xdy/dx + y =dt/dx

or  xdy/dx + t/x =dt/dx

or  dy/dx = 1/x dt/dx - t/x2

so question become     

1/x dt/dx  -1/x2=et  

or  dt/dx    -x et = 1/x

so I.F= e∫-xdx 

           =e-x2/2

    so  t.e-x2/2   = ∫1/x e-x2/2   dx   +c

now it is simple intigration

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin


Tapasranjan Das
14 years ago

Hi Deepak,

dy/dx=exey

=>dy/ey=exdx

=>e-ydy=exdx

=>∫e-ydy=∫exdx

=>-e-y=ex+C

Deepak Aggarwal
14 years ago

Badiuddin sir, actually see the 5th step after u have written the question became. there i think u missed a 't'. actually there the question become

1/x dt/dx  - t/x2=et 

Please check this one out.

Thnks anyways.